OpenStudy (anonymous):

How many units apart are the points (8, 11) and (-1, -1)?

OpenStudy (anonymous):

@andriod09

OpenStudy (anonymous):

to find the distance (or units) apart between two points \[\large (x_1,y_1)~~and~~(x_2,y_2)\] \[\large D=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] D is the distance (or units ) apart between the two points.

OpenStudy (anonymous):

I don't remember how to do any of this

OpenStudy (anonymous):

i have given you the formula.......can you do it now?

OpenStudy (anonymous):

d = √(8- -1)^2 + (11 -1)^2 d = √9^2 + 10^2

OpenStudy (anonymous):

it is supposed to be (11- -1)^2

OpenStudy (anonymous):

after you have corrected the mistake, what do you get??

OpenStudy (anonymous):

d = √9^2 + 12^2 d= √81 + 144 d = √225 d = 15?

OpenStudy (anonymous):

yes.... .....same as i got

OpenStudy (andriod09):

You could of just counted on a graph how far apart they were. LOL its not that hard.

OpenStudy (anonymous):

@andriod09......was that payback for the burn i gave you earlier.. if so, well played

OpenStudy (andriod09):

Huh? I was away had to do the dishes, what burn was this you speak of?

OpenStudy (anonymous):

On one of the questions you helped me with earlier

OpenStudy (anonymous):

that one, remember?